2(a+3a)-78-3b= -14

Answer by richwmiller(17219)

2(a+3a)-78-3b= -14
combine
2*(4a)-78-3b=-14
8a-78-3b=-14
add 78 to both sides
8a-3b=64
This a linear equation which produces a straight line.
You didn't say what you want to do with it.
We can solve for the intercepts when a=o and b=0
We can graph it.
We can get the slope